OpenStudy (homeworksucks):

A 250-g object hangs from a spring and oscillates with an amplitude of 5.42 cm. If the spring constant is 48.0 N/m, determine the acceleration of the object when the displacement is 4.27 cm [down].

OpenStudy (homeworksucks):

I did that and got an acceleration of ~8,19 m/s^2

OpenStudy (homeworksucks):

*8.19

OpenStudy (anonymous):

\[F=ma ; F=kx\] Set those equal to each other to find the acceleration

OpenStudy (homeworksucks):

My only problem is how to find the velocity, I got a max velocity of 0.75 m/s which sounds weird given the acceleration

OpenStudy (homeworksucks):

I used the formula \[v=A \sqrt{\frac{k}{m}}\]

OpenStudy (loser66):

I want to confirm something. Is it ODE for oscillation harmonic function or physics?

OpenStudy (homeworksucks):

Physics, I posted it to Physics, but didn't have any luck

OpenStudy (loser66):

ok, @ybarrap It's physics, friend, help him please

OpenStudy (loser66):

we have \(\omega^2 = \dfrac {k}{m}= \dfrac{48.0}{0.25}= 192\) and acceleration \(\dfrac{d^2x}{dt^2}= -\omega^2 A cos (\omega t+ \phi)\)

OpenStudy (loser66):

@ybarrap is here, wait for him, hehehe. you will have a neat explanation.

OpenStudy (ybarrap):

 $$ x(t)=A\cos(\omega t + \phi)\\ A=.0542\text{ m}\\ \text{Acceleration = }\cfrac{d^2x(t)}{dt^2}=-A\omega^2\cos(\omega t + \phi)\qquad(\star) $$ \(x(0)=0\), so \(\phi=\cfrac{\pi}{2}\) We need to know when \(x(t)=.0427m\). So, $$ x(t)=.0427=-.0542\times\cfrac{48}{.250}\cos(\sqrt{\cfrac{48}{.250}} t + \cfrac{\pi}{2}) $$ After you solve for \(t\), insert back into \(\star\) above and you are done. Make sense?